In the equilibrium SO2Cl2(g) - SO2(g) +Cl2(g) at 2000K and 10 atm pressure, %Cl2 = %SO2 = 40 (by volume)then calculate Kp ?

Grade:12th pass

2 Answers

Arun
5 years ago
Dear student
 

SO2Cl2 = SO2 + Cl2

100               0           0

20                  40       40

2 atm             4 atm   4 atm

now calculate Kp

 

you will get answer as Kp = 8

pranav
4 years ago
volume percent mean given mole percent thats mean it is mole fraction 
 
cl2 mole fraction is then =.4 
then partial pressure =mole fraction  *  total pressure 
=.4*10=4 
same for so2
 
for so2cl2 at eq P - 4 becoz we take so2cl2 initial partial pressure p 
 
now P -4+4+4 should equal to 10
here P =6 
then at eq P -4 =2 
 
now kp =4*4/2 =8